OpenStudy (anonymous):

hello i need help with this Dif eq 1) dy/dx= x2 + y-1

OpenStudy (precal):

What are you doing to your differential equation?

OpenStudy (phi):

Are you studying equations of the form: dy/dx + P(x) y = Q(x)

OpenStudy (stormfire1):

\[dy/dx = x^2 + y -1\]\[dy/dx-y= x^2-1\]\[dy-y=(x^2-1)dx\]Integrate both sides to get:\[-y^2/2=(x^3-3x)/3\]

OpenStudy (stormfire1):

+ C :)

OpenStudy (phi):

@storm clearing the dx will give you a -y dx on the LHS

OpenStudy (stormfire1):

@phi: Good catch...I should have typed in dy(-y) on the LHS...or put -ydy (simpler). For the original poster's info, the final answer is still correct

OpenStudy (phi):

No, your equation becomes \( dy - y dx = (x^{2}-1) dx \) and you can not separate the x and y. See http://www.wolframalpha.com/input/?i=solve+dy%2Fdx+%3D+x%5E2+%2B+y+-1